(sin x / cos 3x)+(sin 3x /cos 9x)+(sin 9x/cos 27x)=(1/2)*(tan 27x- tan x)

Grade:11

3 Answers

suryakanth AskiitiansExpert-IITB
13 years ago

Dear Divya,

Lets consider (sinx / cos3x)

(sinx / cos3x)
= (1/2) * [(2 * sinx * cosx) / (cosx * cos3x)]
= (1/2) * [(sin2x) / (cosx * cos3x)]
= (1/2) * [sin(3x - x) / (cosx * cos3x)]
= (1/2) * [(sin3x * cosx - cos3x * sinx) / (cosx * cos3x)]
= (1/2) * [{(sin3x * cosx) / (cosx * cos3x)} - {(cos3x * sinx) / (cosx * cos3x)}]
= (1/2) * [(sin3x/cos3x) - (sinx/cosx)]
= (1/2) * (tan3x - tanx)

so, (sinx / cos3x) = (1/2) * (tan3x - tanx) ---------(1)

Now putting 3A in place of A in equation (1) we get
(sin3x / cos9x) = (1/2) * (tan9x - tan3x) ------------(2)

Again putting 9A in place of A in equation (1) we get
(sin9x / cos27x) = (1/2) * (tan27x - tan9x) ------------(3)

Now doing (1)+(2)+(3) we get
(sinx / cos3x) + (sin3x / cos9x) + (sin9x / cos27x)
= (1/2) * (tan3x - tanx + tan9x - tan3x + tan27x - tan9x)
= (1/2) * (tan27x - tanx)

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum..

sky
6 years ago
16yrfhu5tthu ucuc7f7f7f77c8g9vb0npnovyxuv8c6uc8g9g 8g8gg87ff78gv99v9b0h0jb099vc87cc78c9vn00jguuc8v8v8v8v8v8vv99h9b9b0j0jn09vv98cc77c7c7c7c7c
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your problem below.
 
Let us consider (sinx / cos3x)
(sinx / cos3x)
= (1/2) * [(2 * sinx * cosx) / (cosx * cos3x)]
= (1/2) * [(sin2x) / (cosx * cos3x)]
= (1/2) * [sin(3x - x) / (cosx * cos3x)]
= (1/2) * [(sin3x * cosx - cos3x * sinx) / (cosx * cos3x)]
= (1/2) * [{(sin3x * cosx) / (cosx * cos3x)} - {(cos3x * sinx) / (cosx * cos3x)}]
= (1/2) * [(sin3x/cos3x) - (sinx/cosx)]
= (1/2) * (tan3x - tanx)
 
So, (sinx / cos3x) = (1/2) * (tan3x - tanx) ---------(1)
Now putting 3A in place of A in equation (1) we get
(sin3x / cos9x) = (1/2) * (tan9x - tan3x) ------------(2)
Again putting 9A in place of A in equation (1) we get
(sin9x / cos27x) = (1/2) * (tan27x - tan9x) ------------(3)
 
Now adding equation (1)+(2)+(3) we get
(sinx / cos3x) + (sin3x / cos9x) + (sin9x / cos27x)
= (1/2) * (tan3x - tanx + tan9x - tan3x + tan27x - tan9x)
= (1/2) * (tan27x - tanx)
= RHS
LHS = RHS
Hence proved.
 
Thanks and regards,
Kushagra